Igloo Industries just paid a dividend of $1.20 per share. The dividends are expected to grow at a constant rate of 5% indefinitely. If investors require a return of 13% on Indigo shares, what is the current price? What should the price be in 4 years time? Plz, list the formula
Answer #1

Calculation of current price:

D0= 1.20

D1= D0*(1+growth)= 1.20*(1+0.05)= 1.26

Growth= 5%

Current price= D1/(return-growth)

Current price= 1.26/(0.13-0.05)= 1.26/0.08= $15.75

Current price is $15.75

Price in 4 years= current price*(1+growth)^4

Price in 4 years= 15.75*(1+0.05)^4= 19.14

Price in 4 years is $19.14